OpenStudy (anonymous):

y=1/2x^2-6x-3 what is the vertex/how to find.., what is the axis of symmetry, range and exact value for x intercepts?

OpenStudy (cwrw238):

= 1/2(x^2 - 12 x - 6) = 1/2[x - 6)^2 - 36 - 6)] = 1/2((x-6)^3 - 42) the vertex is at x = -6 , y = -42* 1/2 = 21 = ( -6,21)

OpenStudy (anonymous):

?? i dont get what u did.. oh u comple ted the square?

OpenStudy (cwrw238):

the x - intercepts occur when y = 0 x^2 - 12x - 6 = 0

OpenStudy (anonymous):

how did u get -6,21?

OpenStudy (cwrw238):

yes - sorry - i should have explained - i did complete the square this changed the function to the vertex form - the x cood. of the vertex is the -6 in the parentheses

OpenStudy (cwrw238):

1/2 * 42

OpenStudy (cwrw238):

3rd line should be (x - 6)^2 not (x-6)^3

OpenStudy (anonymous):

completing the square is stumping me.. it's not adding up to yours for spme reason..

OpenStudy (anonymous):

shouldnt it be +36 since -12 halved is -6 squared is positive 36?

OpenStudy (cwrw238):

go we have x^2 - 12x - 6 divide -12 by 2 to give -6 = (x - 6)^2 - 6 the expansion of x-6 gives x^2 - 12 + 36 so we need -36 and the -6 gives us the -42

OpenStudy (anonymous):

ok i understand vertex.. how do i get the axis of symmetry and range?

OpenStudy (cwrw238):

the range of the function is >= the lowest value which is -21 range ix [-21,+infinity]

OpenStudy (anonymous):

y>-21

OpenStudy (cwrw238):

the axis of symmetry is the vertical line thru the vertex which is x = 6 i'm sorry - its early in morning and i made a mistake with vertex (x-6)^3 give x = +6 as the vertex vertex is at (6,-21) NOT (-6,21)